# Early Jamming Detection in Mobile Indoor Scenarios via Deep Learning

### v1

@@ V-D Impact of Different Hardware and Sampling Rate
We can distinguish two effects from the reported results. Taking into account the configuration with $ROR=1$ and $JOR=1$, we notice that the solution in [7] already reports very low TPR ($0.03$). We highlight that this result is not related to the usage of different oversampling rates but to the new hardware used for the experiments. Indeed, the LimeSDR is a cheaper hardware, which introduces additional inaccuracies and imperfections in the I-Q samples received. Such inaccuracies affect the shape of the I-Q samples, which is now more spread around the expected symbol than before, leading to jammed images different from the ones in the training set of the solution in [7]. These variations are evidently not caught by the approach in [7], which fails to perform reliable jamming detection. On the contrary, in such a configuration, the performance of BloodHound$+$ does not change compared to the results shown in Section V-C, demonstrating once again the enhanced robustness offered by the autoencoders used in BloodHound$+$ compared to the CNNs used in the competing solution.
With higher values of ROR and JOR, the performance of the approach in [7] remains well below $0.2$, confirming the unsuitability of such a solution for the detection of jammers in the wild. Instead, BloodHound$+$ can mitigate and overcome the impact on the wireless channel of different oversampling ratios, being able to detect jamming also when the JOR is very high.
V-E Deceptive Jamming
@@
In previous experiments, we mainly considered two jamming models, that is, tone jamming (using a sinusoid signal) and random jamming (using AWGN). When the adversary does not know the modulation used by the legitimate communication link, the jamming models mentioned and investigated above are the most reasonable options to disrupt the channel as much as possible. However, more powerful attackers might know in advance or become aware at runtime (e.g., by eavesdropping) of the modulation used by the legitimate communication link. Based on such knowledge, they might use the signal as part of a jamming attack optimized to disrupt the communication link as much as possible. Recall that in our experiments the legitimate communication link adopts the BPSK modulation scheme. In such a scenario, the usage of a jamming signal modulated also as a BPSK allows to boost the effectiveness of the jamming activity, as acknowledged by several scientific contributions both analytically [23] and experimentally [24].
@@
To this end, using the same setup as the previous experiments in Section V-D, we evaluated the ability of both BloodHound$+$and the solution in [7] to detect deceptive jamming, i.e., a jammer injecting the same signal used as part of the legitimate communication link. Fig. 14 summarizes the results of our analysis in terms of TPR (for comparison purposes). Note that, also in this case, the TNR of BloodHound$+$ is 0.987, on average over all JOR, since we trained the autoencoder in BloodHound$+$ with the same data as in Fig. 13.
@@
Figure 14: True Positive Ratio (TPR) of BloodHound$+$ and the proposal in [7], with various ROR and JOR values, when the jammer injects the same signal of the legitimate communication link (deceptive jamming).
@@
Even when the adversary uses deceptive jamming, BloodHound$+$ significantly outperforms the benchmark solution in all tested configurations, showing perfect TPR and being also robust with respect to the adoption of a high JOR by the adversary.
@@
Overall, the results reported above demonstrate the superiority of BloodHound$+$ compared to benchmark approaches, as well as the robustness of our solution to a wide range of configuration parameters and scenarios, making it the preferred solution for early jamming detection.
VI Related Work
In the scientific community, jamming detection is usually achieved by applying various types of analysis on one or more metrics extracted from the primary communication link. Regarding the metrics, several parameters have been analyzed, such as the RSS of the signals [29], the PDR [30] as in [31], the profile of packet retransmissions, as in [32], or modulation-specific metrics, such as for Orthogonal Frequency Division Multiplexing (OFDM) in [33].
At the same time, due to the increasing popularity of Artificial Intelligence (AI), Machine Learning (ML) and DL approaches have been recently used extensively for detecting ongoing jamming. Such tools include CNNs such as in [34] and [35], genetic algorithm-based cumulative sum methods such as in [36], Bayesian networks such as in [37] and, finally, autoencoders such as in [38]. All such approaches utilize as the main source of information the PHY layer of the communication stack, due to its direct relationship with the wireless channel, where jamming occurs. However, although some of the cited contributions analyzed the performance of the proposed jamming detection technique with low SNR, none of them considered the BER of the communication link. As a result, the proposed approaches mostly confirm that the root cause of the drop in BER is jamming. However, they cannot detect such attacks even when the jamming effect is so low as not to affect the BER of the communication link. As explained above, such consideration is particularly relevant in mobile scenarios for remotely controlled equipment, not to lose control of the mobile entity completely before detecting jamming. In this context, the only contribution in the literature achieving such a property is our previous proposal in [7]. As shown in Section V the methodology shown in this paper significantly outperforms both the one proposed in [7] and the evolutions of such a methodology as the one proposed in [8].
@@ VII Conclusion
In this paper, we have presented BloodHound$+$, an approach that allows drones and possibly other mobile devices to detect jamming at the physical layer (PHY) of the communication stack. Our solution works on raw I-Q samples extracted from the communication link, converts them into grayscale images, and uses sparse autoencoders to detect discrepancies with the expected profile of the channel. Therefore, BloodHound$+$ can detect jamming well before its effect could cause a significant decrease in the quality of the main communication link. In turn, our solution allows drones to efficiently avoid the jammed area and maintain full control and safety. To test the effectiveness of our solution, we conducted an extensive measurement campaign, acquiring real-world data with different hardware, jamming strategies, and scenario configurations. We also tested the performance of BloodHound$+$ depending on various parameters, such as SNR of the communication link, the distance from the jammer and the transmitter, the size of the training set, the number of samples acquired from the channel, the jammer oversampling ratio, and the jamming strategies. Our experimental assessment demonstrates, through a variety of results, the superiority of our solution compared to the current state-of-the-art across all the analyzed configuration parameters.
In general, we believe our solution contributes to taking a step further toward the safe and secure integration of drones into daily life. As part of our future work, we plan to further investigate the effectiveness of BloodHound$+$ when applied for outdoor applications.
@@ Acknowledgements
This work has been partially supported by the INTERSECT project, Grant No. NWA.1162.18.301, funded by the Netherlands Organization for Scientific Research (NWO). Any opinions, findings, conclusions, or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of NWO. Moreover, this research was sponsored in part by the NATO Science for Peace and Security Programme under grant SPS G5797.

### v4

@@ VI-D Impact of Different Hardware and Sampling Rate
We can distinguish two effects from the reported results. Considering the configuration with $ROR=1$ and $JOR=1$, we notice that the solution in [16] already reports very low TPR ($0.03$). We highlight that this result is not related to different oversampling rates but to the new hardware used for the experiments. Indeed, the LimeSDR is a cheaper hardware, which introduces additional inaccuracies and imperfections in the I-Q samples received. Such inaccuracies affect the shape of the I-Q samples, which is now more spread around the expected symbol than before, leading to jammed images different from the ones in the training set of the solution in [16]. The approach in [16] does not catch these variations, thus failing to perform reliable jamming detection. On the contrary, in such a configuration, the performance of BloodHound$+$ does not change compared to the results shown in Section VI-C, demonstrating once again the enhanced robustness offered by the autoencoders used in BloodHound$+$ compared to the CNNs used in the competing solution. With higher ROR and JOR values, the performance of the approach in [16] remains well below $0.2$, confirming the unsuitability of such a solution for detecting jammers in the wild. Instead, BloodHound$+$ can mitigate and overcome the impact on the wireless channel of different oversampling ratios, being able to detect jamming also when the JOR is very high.
VI-E Deceptive Jamming
@@
In previous experiments, we mainly considered two jamming models, that is, tone jamming (using a sinusoid signal) and random jamming (using AWGN). When the adversary does not know the modulation used by the legitimate communication link, the jamming models mentioned and investigated above are the most reasonable options to disrupt the channel as much as possible. However, more powerful attackers might know in advance or become aware at runtime (e.g., by eavesdropping) of the modulation used by the legitimate communication link. Based on this knowledge, they might use the signal as part of a jamming attack optimized to disrupt the communication link as much as possible. Recall that the legitimate communication link in our experiments adopts the BPSK modulation scheme. In such a scenario, the usage of a jamming signal modulated also as a BPSK allows to boost the effectiveness of the jamming activity, as acknowledged by several scientific contributions both analytically [58] and experimentally [28].
@@
To this end, using the same setup as the previous experiments in Section VI-D, we evaluated the ability of both BloodHound$+$ and the solution in [16] to detect deceptive jamming, i.e., a jammer injecting the same signal used as part of the legitimate communication link. Fig. 15 summarizes the results of our analysis in terms of TPR (for comparison purposes). Note that also in this case, the TNR of BloodHound$+$ is 0.987, on average over all JOR, since we trained the autoencoder in BloodHound$+$ with the same data as in Fig. 14.
@@
Figure 15: True Positive Ratio (TPR) of BloodHound$+$ and the proposal in [16], with various ROR and JOR values, when the jammer injects the same signal from the legitimate communication link (deceptive jamming).
@@
Even when the adversary uses deceptive jamming, BloodHound$+$ significantly outperforms the benchmark solution in all tested configurations, showing perfect TPR and being robust to adopting a high JOR by the adversary.
@@
Overall, the results reported above demonstrate the superiority of BloodHound$+$ compared to benchmark approaches and the robustness of our solution to a wide range of configuration parameters and scenarios, making it the preferred solution for jamming detection in a low-BER regime.
VI-F Considerations on Interference
Various factors, such as interference and obstacles, might affect the performance of BloodHound$+$. We show in Fig. 16 the effect on the I-Q samples of people passing through our experimental setup.
(a) t = 0
(b) t = 1
(c) t = 2
(d) t = 3
(a) t = 0
(b) t = 1
(c) t = 2
(d) t = 3
(e) t = 4
Figure 16: The effect of multipath on the transmitter-receiver link across five time windows. Compared to an ideal (static) scenario (Figs. (a) and (e)), the shape of the I-Q samples at the receiver (modulated through the BPSK scheme) is significantly affected (see Figs. (b), (c), and (d)) when an event is happening, e.g., moving objects.
We notice that the shape of the I-Q samples at the receiver is significantly affected compared to a LOS scenario. Also, compared to jamming, note that these phenomena are much quicker, and the shape of the I-Q samples returns to the regular shape much faster than under jamming attacks. Thus, to allow BloodHound$+$ not to declare jamming in such cases, we need to experience such phenomena during the training phase to make them part of the regular conditions of the communication channel, especially indoors. These considerations motivate the deployment of our solution and the experiments in an office environment during working hours: we experience several instances of such phenomena during training, and they contribute to creating an expected profile of unjammed scenarios which takes interferences into account. During jamming attacks, the shape of the I-Q samples is displaced from the expected profile for a much extended time, contributing to enhancing the performance of our solution.
@@ VII Conclusion
In this paper, we have presented BloodHound$+$, an approach that allows drones and possibly other mobile devices to detect jamming at the physical layer (PHY) of the communication stack. Our solution works on raw I-Q samples extracted from the communication link, converts them into grayscale images, and uses sparse autoencoders to detect discrepancies with the expected profile of the channel. Therefore, BloodHound$+$ can detect jamming in low-BER regimes, i.e., well before its effect could cause a significant decrease in the quality of the main communication link. At the same time, our solution allows drones to efficiently avoid the jammed area and maintain complete control and safety. To test the effectiveness of our solution, we conducted an extensive measurement campaign, acquiring real-world data with different hardware, jamming strategies, and scenario configurations. We also tested the performance of BloodHound$+$ depending on various parameters, such as SNR of the communication link, the distance from the jammer and the transmitter, the size of the training set, the number of samples acquired from the channel, the jammer oversampling ratio, and the jamming strategies. Our experimental assessment demonstrates, through an extensive collection of results, the superiority of our solution compared to the current state-of-the-art across all the analyzed configuration parameters. In general, our solution contributes to taking a step further toward the safe and secure integration of drones into daily life. As part of our future work, we plan to investigate further the effectiveness of BloodHound$+$  e.g., when applied for outdoor applications.
@@ Acknowledgements
This work has been partially supported by the INTERSECT project, Grant No. NWA.1162.18.301, funded by the Netherlands Organization for Scientific Research (NWO). Any opinions, findings, conclusions, or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of NWO. Moreover, this publication was made possible by the NPRP12C-0814-190012-SP165 awards from the Qatar National Research Fund (a member of Qatar Foundation).